OpenStudy (suwhitney):

You own a 10 year 6% simple interest bearing note with a face value (principal) of $10,000. If you decide to sell the note two years early and receive a sale price of $14,000 for the note, what simple interest rate did you actually earn for the 8 years that you owned the note?You earned ______ percent. Give answer to the nearest whole percent.Give numeric value only without including the percent sign in your answer.

OpenStudy (anonymous):

A=p(r)(t)+p so A=10,000(.05)(8)+10,000 A=14,000 so ur answer would be 5%

OpenStudy (suwhitney):

Thank you so much!! :) That's easier than it seems :P

OpenStudy (anonymous):

No problem :)